How do I sketch this graph? Please explain how to.
- 11 DUO ........ 10006 9000 3000...... 1000 sou 4000 zo0 Qi sketch..graph...of y= 10.500 - 1, 500x.
Answer #1

given Y = 10500 -- 1500m find a when yco 0 = 10500 -1500 % 150044 =10500 - 10500 25 T500 18= y when 720 10500 - 1500 to y = 1